# SwiftBrush[Uncaptioned image]: One-Step Text-to-Image Diffusion Model with Variational Score Distillation

### v1

@@ 4.3 Analysis
The importance of student parameterization. Apart from the LoRA teacher, the re-parameterization of the student (Sec. 3) also plays an important role in the SwiftBrush training. Without the it, the student is unable to to create realistic images and it only produces noisy results, as can be noticed from the $4^{th}$ column of Fig. 5. This can be further backed up by the orange curve in both FID and CLIP score in Fig. 4 (a) and (b), where student without re-parameterization perform consistently worst among all models.
@@
Interpolating Latent. We conduct a controlled experiment where the noise input $z$ remains the same while varying the input prompt $y$ (Fig. 6). Interestingly, the SwiftBrush-distilled student seemingly features enhanced control and editability compared to the original diffusion model due to its one-step nature. Additionally, we present further results of latent space interpolation in Sec. 7. In this process, the student model, distilled from the pretrained Stable Diffusion teacher, demonstrates a seamless transition in its image generation.
5 Conclusion and Future Work
In this work, we have introduced a new image-free distillation scheme, SwiftBrush, which leverages insights from text-to-3D synthesis to speed up text-to-image generation. Empirically, the results presented in this paper demonstrate the effectiveness of our approach in making text-to-image generation more efficient and accessible. We hope our work will encourage broader use of text-to-image generation systems by substantially lowering the inference cost.
SwiftBrush inevitably produce lower quality samples compared to the teacher which utilizes multi-step sampling for inference. Another drawback is that the current design only focuses on single-step student model and cannot support few-step generation as did in Luo et al. [22] to improve image fidelity. In future work, we aim to investigate the feasibility of extending SwiftBrush to a few-step generation so that we can trade computation for quality. Furthermore, we also find it compelling to explore SwiftBrush-like training where only one teacher is required. Such an exploration could provide tremendous benefits in scenarios where the computational resource is limited. Finally, it is intriguing to see if techniques such as DreamBooth [33], ControlNet [47] or InstructPix2Pix[2] can be integrated with SwiftBrush, allowing instant generation of various applications.
@@ 5 Conclusion and Future Work
6 Training Cost and Inference Speed
@@
Inference Speed. Next, we measure the inference time of our method using a batch size of $1$ on an A100 40GB GPU. The inference time of one-step UNet to generate the latent from random noise is $25$ ms, while those of text encoder and VAE are $10$ ms and $5$ ms, respectively. Consequently, the total inference time of SwiftBrush is $25+10+5=40$ ms. Furthermore, the speed of the text encoder and VAE in BOOT are the same as ours; however, there is a small overhead in the UNet ($30$ms) as additional layers are used according to their implementation guidelines. Therefore, the total inference time for BOOT is $45$ms.
Compare with Others. We also sum up inference speed and training time of existing methods and ours in Tab. 3. Since Guided Distillation [23] does not provide details on the inference time or access to the source code, we have decided not to include its speed in our report. Also, we use the provided pretrained model to measure the inference time of LCM [22] which is slightly faster than ours. In terms of training costs, Guided Distillation and InstaFlow [19] require a considerable amount of time, limiting their practical use. Conversely, LCM converges quickly but produces poor-quality images with one-step inference. Our method, however, offers both training efficiency and high-quality output.
Method Inference Time (ms)$\downarrow$ Training Time (days) $\downarrow$
Guided Distillation - ${}^{\dagger}$
[…]
Instaflow ${}^{\dagger}$ ${}^{\dagger}$
BOOT ${}^{\ddagger}$ ${}^{\ddagger}$
Ours 40 4.11
[…]
@@ 7 Additional Visual Results
Uncurated Samples. We display uncurated images conditioned on 25 random text prompts. The images generated by various models are presented in Fig. 7, 8, 9 and 10, allowing for better visual comparison between our model and other methods.
Additional Visuals of Analysis. To further back up our claim in Sec. 4.3, we provide some uncurated samples using the text prompt “An DSLR photo of a tiger in the city“. The images generated by four configurations at every 2000 iterations are presented in Fig. 13, allowing for a better visual comparison of each component’s importance in our proposed method across different training stages.
Additional Visuals of Interpolation. Apart from investigating the role of each component in SwiftBrush, we also explore its properties when interpolating text prompt condition $y$ (Fig. 11) or noise input $z$ (Fig. 12). It is evident that when the input text prompt $y$ is interpolated, our model generates a seamless transition in the output, showcasing excellent editability and controllability. On the other hand, it is noteworthy that the quality is considerably lower when interpolating the noise input $z$, leading to a poor transition between two endpoints.
@@
Figure 11: Results of interpolating the input prompt. The prompts used here are selected from a standard template: “A DSLR photo of a {animal} reading a book”. Here, ‘animal’ is dog or cat and we interpolate the text embedding using alpha-blending. Same noise input $z$ is used for images at each row.
Figure 12: Results of interpolating the noise input. The prompts used here are selected from a standard template: “A DSLR photo of a {animal} reading a book”. Here, ‘animal’ is dog, cat, owl or panda and we interpolate the noise input using alpha-blending. Same text input $y$ is used for images at each row.

### v4

@@ 4.3 Analysis
The importance of student parameterization. Apart from the LoRA teacher, the re-parameterization of the student (Sec. 3) also plays an important role in the SwiftBrush training. Without it, the student is unable to to create realistic images and it only produces noisy results, as can be noticed from the $4^{th}$ column of Fig. 5. This can be further backed up by the orange curve in both FID and CLIP score in Fig. 4 (a) and (b), where student without re-parameterization perform consistently worst among all configurations.
@@
Interpolating Latent. We conduct a controlled experiment where the noise input $z$ remains the same while varying the input prompt $y$ (Fig. 6). Interestingly, the SwiftBrush-distilled student seemingly features enhanced control and editability compared to the teacher model due to its one-step nature. Additionally, we present further results of latent space interpolation in Sec. 7. In this process, the student model, distilled from the Stable Diffusion teacher, demonstrates a seamless transition in its image generation.
5 Conclusion and Discussion
Conclusion. In this work, we have introduced a new image-free distillation scheme, SwiftBrush, which leverages insights from text-to-3D synthesis to speed up text-to-image generation. Empirically, the results presented in this paper demonstrate the effectiveness of our approach in making text-to-image generation more efficient and accessible. We hope our work will encourage broader use of text-to-image generation systems by greatly lowering the inference cost.
Discussion. SwiftBrush inevitably produces lower quality samples compared to the teacher which utilizes multi-step sampling for inference. Another drawback is that the current design only focuses on single-step student model and cannot support few-step generation as did in Luo et al. [23] to improve image fidelity. In future work, we aim to investigate the feasibility of extending SwiftBrush to a few-step generation so that we can trade computation for quality. Furthermore, we also find it compelling to explore SwiftBrush-like training where only one teacher is required. Such an exploration could provide tremendous benefits in scenarios where the computational resource for training is limited. Finally, it is intriguing to see if techniques such as DreamBooth [35], ControlNet [49] or InstructPix2Pix[2] can be integrated with SwiftBrush, allowing instant generation of various applications.
@@ 5 Conclusion and Discussion
6 Additional Details
@@
Inference Speed. Next, we measure the inference time of our method using a batch size of $1$ on an A100 40GB GPU. The inference time of one-step UNet to generate the latent from random noise is $25$ ms, while those of text encoder and VAE are $80$ ms and $5$ ms, respectively. Consequently, the total inference time of SwiftBrush is $25+80+5=110$ ms. Furthermore, the speed of the text encoder and VAE in BOOT are the same as ours; however, there is a small overhead in the UNet ($30$ms) as additional layers are used. Therefore, the total inference time for BOOT is $115$ms.
Memory Requirement. Furthermore, we evaluate SwiftBrush’s memory consumption during both training (with a batch size of 16) and inference (with a batch size of 1) on an A100 GPU. Since our approach involves training only a student and an additional low-rank LoRA weight for our LoRA teacher, it maintains a minimal memory footprint.
Compare with Others. We also sum up inference speed, training time and memory requirements of existing methods and ours in Tab. 3. In terms of training costs, Guided Distillation and InstaFlow [20] require a considerable amount of training time, limiting their practical use. Conversely, LCM converges quickly but produces poor-quality images with one-step inference. Our method, however, offers both efficiency and high-quality output.
Method Training Inference
Time (GPU days) Memory (GB) Time (ms) Memory (GB)
Guided Distillation 108.8 - - -
LCM 1.3 33.6 118 11.3
InstaFlow 199.2 - 116 12.3
BOOT 5.6 30.2 115 8.3
Ours 4.1 26.4 110 8.2
Table 3: Comparison of our method against other works on inference and training time/memory. means that we obtain the numbers from the corresponding papers. means that we obtain the numbers using the official code of the corresponding papers. means that we re-implement the work and report the numbers. The units for inference speed, training time and memory usage are miliseconds, A100 GPU days and gigabytes, respectively.
@@ 7 Additional Visual Results
Uncurated Samples. We display uncurated images conditioned on 25 random text prompts***https://thuanz123.github.io/swiftbrush/assets/prompts.txt. The images generated by various models are presented in Fig. 7, 8, 9 and 10 for a better visual comparison between our model and others.
Additional Visuals of Analysis. To further back up our claim in Sec. 4.3, we provide some uncurated samples using the text prompt “An DSLR photo of a tiger in the city“. The images generated by four configurations at every 2000 iterations are presented in Fig. 13 for a better visual comparison of each component’s importance in our proposed method across different training stages.
Additional Visuals of Interpolation. Apart from investigating the role of each component in SwiftBrush, we also explore its properties when interpolating text prompt condition $y$ (Fig. 11) or noise input $z$ (Fig. 12). It is evident that our model generates a seamless transition in the output, showcasing excellent editability and controllability.
8 Additional Quantitative Results
We run SwiftBrush on CIFAR-10 $32\times 32$ and class-conditional ImageNet $64\times 64$ benchmarks using EDM teacher models and compare it with Progressive Distillation (PD) and Consistency Distillation (CD), either using L2 or LPIPS loss, in Tab. 4. SwiftBrush is on par with CD-LPIPS and it significantly outperforms the others.
@@ 8 Additional Quantitative Results
Figure 11: Results of interpolating the input prompt. The prompts used here are selected from a standard template: “A DSLR photo of a {animal} reading a book”. Here, ‘animal’ is dog or cat and we interpolate the text embedding using linear interpolation (Lerp). Same noise input $z$ is used for images at each row.
Figure 12: Results of interpolating the noise input. The prompts used here are selected from a standard template: “A DSLR photo of a {animal} reading a book”. Here, ‘animal’ is dog, cat, owl or racoon and we interpolate the noise input using spherical linear interpolation (Slerp). Same text input $y$ is used for images at each row.
@@
Method Steps CIFAR-10 $32\times 32$ ImageNet $64\times 64$
PD 1 8.34 15.39
CD - L2 1 7.90 12.10
CD - LPIPS 1 3.55 6.20
Ours 1 4.46 5.85
EDM 35 1.97 -
EDM 79 - 2.44
Table 4: Comparison of our method against other works in FID on CIFAR-10 ($32\times 32$) and ImageNet ($64\times 64$) benchmarks.